# Representative Ranking for Deliberation in the Public Sphere

### v1

@@ Abstract
**footnotetext: Equal contribution
@@ 1 Introduction
In theories of deliberative democracy (Habermas, 1991; Cohen, 2005; Fishkin, 2009; Bächtiger et al., 2018; Landemore, 2020), the legitimacy of collective decisions is grounded in the thoughtful and inclusive public deliberation that produces them. The deliberation that occurs in formal environments (e.g. a parliamentary meeting or citizens’ assembly) can be contrasted to the widespread and informal “deliberation in the wild” that happens in the public sphere (e.g. in coffee shops) (Habermas, 1991, 2023). While the formal environments can be designed to provide ideal conditions for constructive deliberation, they do not easily scale—most people will not have the opportunity to regularly participate in citizen’s assemblies or parliamentary meetings. Therefore, improving informal public sphere deliberation may be essential for improving deliberation overall (Lazar and Manuali, 2024; Landemore, 2024).
@@
Online platforms, such as comment sections on news sites or social media, present an opportunity to expand public sphere deliberation (Helberger, 2019). These digital spaces enable interactions between individuals who might never meet in person, potentially exposing them to a broader array of viewpoints. However, in reality, these online discussions often deteriorate into low-quality or toxic exchanges (Nelson et al., 2021; Kim et al., 2021). In this work, we aim to bring online conversations closer towards the ideals of deliberation, leveraging algorithmic tools in doing so. While we do not expect that comment sections will achieve deliberative ideals to the same extent that in-person formal environments such as citizen assemblies do, we believe that striving towards deliberative ideals can be a valuable guiding principle for improving algorithmic interventions for these spaces.
@@
What are the ideals that make a good deliberation? Deliberative democracy ideals generally demand at minimum (i) certain aspects of conversational quality (e.g. respect, reasoned arguments, etc), and (ii) the representation of diverse voices. For example, Fishkin and Luskin (2005) claim that deliberative discussion must be (i) conscientious, i.e., “the participants should be willing to talk and listen with civility and respect,” and (ii) comprehensive, i.e., “all points of view held by significant portions of the population should receive attention.” Similarly, in their review of the ‘second generation of deliberative ideals,’  Bächtiger et al. (2018) highlight the ongoing importance of (i) respect, and (ii) the evolving view of equality as the equality of opportunity to political influence (Knight and Johnson, 1997), which requires that people with diverse viewpoints be able to receive attention to their perspectives.
@@
Algorithmic moderation and ranking have become key tools for facilitating conversational quality in online discussions (Kolhatkar et al., 2020). For example, Google Jigsaw’s Perspective API comment classifiers have been used by prominent news organizations, such as the New York Times and Wall Street Journal, to filter or rerank comments to improve civility (Lees et al., 2022; Saltz et al., 2024). More recently, a line of work on bridging systems advocating ranking aims at bridging different perspectives and reducing polarization (Ovadya and Thorburn, 2023a). Bridging-based ranking has been used to in several high-impact applications: ranking comments on Facebook (Ovadya and Thorburn, 2023a), selecting crowd-sourced fact checks on X (Wojcik et al., 2022), and selecting comments in collective response systems like Pol.is and Remesh (Small et al., 2021; Konya et al., 2023a; Huang et al., 2024).
@@
Figure 1: Our approach compared to the standard approach to ranking comments. In this example, a platform with $n=6$ users wants to select $k=3$ comments to highlight while optimizing a given score function $f$ (e.g. engagement, civility, diverse approval, etc). This example is inspired by our experiments ranking comments related to campus protests in Sec 6. In the standard approach, the platform simply selects the $k$ comments with the highest score (in this case, diverse approval across the red and blue users). However, this leads to only two users having a comment that they approve of in the selected set $S^{*}$. The two remaining red and blue users do not receive representation despite the fact that they each form a group large enough to warrant representation by proportionality (each of size $\lceil n/k\rceil\geq\lceil 6/3\rceil=2$) and are minimally cohesive (have at least one item that they agree upon in common). On the other hand, in our approach, the platform picks the $k$ comments that maximize the score while satisfying justified representation, and guarantees that all cohesive groups of size at least $n/k$ are represented in the resulting set $S^{*}_{JR}$.
@@
While these interventions may improve conversational quality, the goal of ensuring representation in online discussions is less studied. In fact, solely focusing on conversational quality might inadvertently result in the suppression of certain groups, in the sense that their comments may be filtered out or do not appear high enough in the comment ranking to gain visibility. For example, civility classifiers have been shown to disproportionately flag comments written in African-American english (Sap et al., 2019) and comments discussing encounters with racism (Lee et al., 2024). In the context of diverse approval, the groups being bridged have a large effect on what content gets shown. Commonly, these groups correspond to political groups like left- or right-leaning users (Wojcik et al., 2022). Therefore, it is plausible that diverse approval could give more visibility to moderate comments, while failing to provide representation of others who hold more ideological views. [^1] Or, imagine an online platform with Celtics and Knicks fans whereby the only posts that get cross attention are those about Kadeem Allen (a former player in both franchises). Bridging-based ranking à la Ovadya and Thorburn (2023a) may overly represent the Kadeem Allen posts, downgrading legitimate interest of the broader fan-bases.
@@
    Theoretical framework. We contribute a theoretical framework in which the platform selects $k$ top-ranked comments by maximizing a score function $f$ (e.g. civility, bridging objectives) subject to a representation constraint known as justified representation (JR) (Aziz et al., 2017) from the social choice literature on approval voting in multi-winner elections.
@@
    Theoretical analysis. We theoretically analyze how the JR constraint affects the platform’s ability to maximize different classes of score functions: in the natural setting, where users’ opinions cluster into a few groups, enforcing representation does not come at a price to other objectives such as civility or user engagement.
@@
    Method and empirical analysis. Leveraging approximation algorithms from social choice theory, we show how to implement a JR constraint for comment ranking. In experiments ranking comments on Remesh [^2] Remesh (Konya et al., 2023a) is a popular collective response system (Ovadya, 2023) used frequently by governments, non-profits, and corporations to elicit opinions from a collective at scale. related to campus protests, we found that enforcing JR significantly enhanced representation without compromising other measures of conversational quality.
@@ 2 Related work
Algorithmic interventions for comment ranking focus primarily on promoting conversational quality  (Kolhatkar and Taboada, 2017; Saltz et al., 2024; Lees et al., 2022; Ovadya and Thorburn, 2023a; Jia et al., 2024; Piccardi et al., 2024), but without guarantees on representation of viewpoints—another key ideal of deliberation. To incorporate such guarantees into these methods, we focus on selecting the set of top $k$ comments, which allows us to leverage notions of proportional representation. While other research on diversity in recommender systems also focuses on selecting optimal sets of items, diversity in this context is typically defined based on the content of items, e.g., choosing videos from a variety of topics (Kunaver and Požrl, 2017; Zhao et al., 2025). In contrast, our goal here is to select a set of comments that ensures representation, thereby focusing on the perspective of users rather than the content diversity of the items.
@@
To formalize representation, we borrow the social choice theoretic concept of justified representation (JR), developed by Aziz et al. (2017) in the context of approval voting in multi-winner elections. Justified representation is an axiom that formalizes a notion of proportionality: it guarantees that every large enough group of users that has shared preferences (approves of at least one comment in common) is allocated at least one of the top comments. While prior work has focused on selecting or generating comments subject to JR or extensions of JR (Halpern et al., 2023; Fish et al., 2024), these works do not consider the optimization of other exogenous objectives (e.g. content-based classifiers for civility or other attributes) that are a key component of real-world comment ranking. Therefore, we study a constrained optimization problem where the platform selects the top comments based on optimizing a score function (e.g. civility, diverse approval, engagement, etc) subject to JR. We focus on JR because it is the most fundamental in the family of axioms, however, it would be interesting to explore extensions of JR (Aziz et al., 2017; Sánchez-Fernández et al., 2017; Skowron et al., 2017; Brill and Peters, 2023; Fish et al., 2024) that afford stronger guarantees of representation in future work.
@@
For the comment ranking setting, in particular, JR affords several advantages over notions of demographic or social representation more commonly used in the algorithmic fairness literature (Barocas et al., 2023; Chasalow and Levy, 2021). We discuss these advantages further in Appendix A, but here, we highlight two benefits. First, unlike most algorithmic fairness methods, JR does not require inference of any demographic or social labels, making it compatible with the privacy and legal constraints of real-world platforms (Holstein et al., 2019; Veale and Binns, 2017). Second, JR automatically adapts to the groups that are relevant in different contexts. For instance, the groups relevant to represent in the discussions about the NYC budget planning process differ from those in a post about a basketball game between the Celtics and the Knicks. On real-world platforms that have a diverse range of content, the fact that JR naturally focuses on the relevant groups is a significant advantage.
@@
The primary theoretical question we analyze is: what is the price of enforcing JR, with respect to the ability to optimize other score functions? For instance, is it the case that the platform must endure much higher levels of toxicity in order to provide representation? The closest related work is by Elkind et al. (2022) who investigate the price of JR with respect to the utilitarian social welfare (the number of selected comments that each user approves of, summed over users). We expand their work by analyzing the price for other classes of score functions and constraints on the user approval matrix. We find that in the common setting when users’ preferences cluster into a few groups, the price of JR is typically negligible, which also relates to Faliszewski et al. (2018)’s work who use clustering as an empirical heuristic to achieve representation.
@@ 3.1 Model
For $m\in\mathbb{N},$ we write $[m]=\{1,\cdots,m\}.$ Let $[m]$ be the set of all comments and $[n]$ be a set of users. The platform must select a subset $S\subseteq[m]$ of $k$ comments to highlight. [^3] Work on diversity in recommender systems also focuses on sets of items, but with the goal of showing “diverse” items where diversity is typically defined based on the content of the items, e.g., showing videos spanning different topics or genres (Kunaver and Požrl, 2017; Zhao et al., 2025). In contrast, our focus is on selecting sets of items (comments) to ensure that the top items provide a degree of proportional representation to users. Each user $u\in[n]$ approves a set $A_{u}\subseteq[m]$ of comments and $\mathcal{A}_{n}=\{A_{1},\cdots,A_{n}\}$ is the profile of approved comments across all users. In practice, a user’s approval set might be defined as the set of comments that they upvote, like or react to. In all, each problem is characterized by a tuple $\mathcal{I}_{m,\mathcal{A}_{n},k}=\langle m,\mathcal{A}_{n},k\rangle$ of $n$ approval sets over $m$ comments, from which $k$ comments are selected.
@@ 3.2.2 Approval Dependent Scoring Rules
A scoring rule $f(i,\mathcal{A}_{n})$ is approval dependent if, for all item $i$ not in any approval set of $\mathcal{A}_{n}$ (that is $i\notin\cup_{u=1}^{n}A_{u}$), then we have $f(i,\mathcal{A}_{n})=0$.
@@
Diverse approval is another example of an approval dependent scoring rule, where the score of a comment reflects the level of approval it receives across diverse groups of users. These groups could be pre-defined, e.g., based on demographic characteristics, or more commonly, learned from the data. Diverse approval is motivated by research finding that, compared to pure engagement, diverse approval tends to correlate with comments that are less toxic, more informative, and higher quality (Ovadya and Thorburn, 2023b). Diverse approval has been used to select and rank user-generated content in several high-impact applications: ranking comments on Facebook (Ovadya and Thorburn, 2023a), selecting crowd-sourced fact checks on X (Wojcik et al., 2022), and selecting comments in collective response systems like Pol.is and Remesh (Small et al., 2021; Konya et al., 2023a; Huang et al., 2024).
@@
Other variants have also been used such as the product of the approval rate across groups (Small et al., 2021; Huang et al., 2024) or a softmax version of minimax diverse approval (Konya et al., 2023a). Note that these examples satisfy a property that is strictly stronger than approval dependency, that of approval monotonicity where an approval monotonic function is an approval dependent function such that $\forall(\mathcal{A}^{\prime}_{n}$, $\mathcal{A}_{n}$) with $A^{\prime}_{u}\succeq_{(i)}A_{u}$ for some $u$ and $A^{\prime}_{v}=A_{v}$ for all $v\in[n]\setminus\{u\},$ then $f(i,\mathcal{A}^{\prime}_{n})\geq f(i,\mathcal{A}_{n}).$
@@ 3.3 Justified Representation
The problem of selecting $k$ items from a set $[m]$ based on individual’s approval ballots is well studied in the context of multi-winner social choice, where the $k$ “items” are candidates to be selected for a committee. [^5] An intuitive approach would be to select the $k$ items with the largest approval score |$|\{u\in[n]\mid i\in A_{u}\}|$. However, this method tends to favor items supported by the majority, potentially excluding minority groups from representation. Let a world with $60$ people who approve the $10$ items in set $A$ and another $40$ people who approve of a distinct set $A^{\prime}$ of $10$ items. If a committee of size $10$ is selected based on approval scores, the winning committee, $A$ would fail to represent $40\%$ of the world. Instead, a committee composed of $6$ items from $A$ and $4$ items from $A^{\prime}$ would respect an intuitive notion of proportionality. In order to ensure fairness, in the sense that all large groups with cohesive preferences receive some amount of representation, Aziz et al. (2017) axiomatically formalized the idea of justified representation. Following Elkind et al. (2022)’s exposition, and Bredereck et al. (2019a)’s notion of justifying sets, assume we are given an instance $\mathcal{I}_{m,\mathcal{A}_{n},k}.$ of $n$ approval sets over $m$ comments, from which the platform must select $k$ comments.
@@
The concept of Justified Representation (JR) follows to ensure that large enough cohesive groups are minimally represented.
@@
The strength of the JR axiom comes from its consideration of all cohesive groups of size at least $n/k$. However, it may seem like a minimal requirement for representation, as it requires only one member of a group to approve of a comment in the chosen set for that group to be represented. It appears as though there could be perverse sets that satisfy JR where most people do not approve of the selected comments. However, research has found that common algorithms used to satisfy JR do not lead to such perverse outcomes (Bredereck et al., 2019a). In our experiments in Section 6, we also find that enforcing JR leads to greater coverage, i.e., more users have a comment they approve of in the selected set. In future work, it would also be interesting to explore strengthened versions of JR (Aziz et al., 2017; Sánchez-Fernández et al., 2017; Brill and Peters, 2023; Fish et al., 2024).

### v2

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
In theories of deliberative democracy (Habermas, 1991; Cohen, 2005; Fishkin, 2009; Bächtiger et al., 2018), thoughtful and constructive public discourse is key for a well-functioning polity. Online platforms, such as comment sections on news sites or social media, present an opportunity to expand this kind of public sphere deliberation (Helberger, 2019; Landemore, 2024). These digital spaces enable interactions between individuals who might never meet in person, potentially exposing them to a broader array of viewpoints. However, in reality, these online discussions often deteriorate into low-quality or toxic exchanges (Nelson et al., 2021; Kim et al., 2021). In this work, we aim to bring online conversations closer towards the ideals of deliberation, leveraging algorithmic tools in doing so.
@@
What does it take for a discussion to count as deliberative? Deliberative democracy ideals generally demand at minimum (i) certain aspects of conversational quality (e.g., respect, reasoned arguments, etc) and (ii) the representation of diverse voices. For example, among other ideals, Fishkin & Luskin (2005) claim that deliberative discussion must be (i) conscientious, i.e., “the participants should be willing to talk and listen with civility and respect,” and (ii) comprehensive, i.e., “all points of view held by significant portions of the population should receive attention.” Similarly, in their review of the ‘second generation of deliberative ideals,’  Bächtiger et al. (2018) highlight the ongoing importance of (i) respect, and (ii) the evolving view of equality as the equality of opportunity to political influence (Knight & Johnson, 1997), which requires that people with diverse viewpoints be able to receive attention to their perspectives.
@@
Algorithmic moderation and ranking have become key tools for facilitating conversational quality in online discussions (Kolhatkar et al., 2020). For example, Google Jigsaw’s Perspective API comment classifiers have been used by prominent news organizations, such as the New York Times and Wall Street Journal, to filter or rerank comments to improve civility (Lees et al., 2022; Saltz et al., 2024). More recently, a line of work on bridging systems advocating ranking aims at bridging different perspectives and reducing polarization (Ovadya & Thorburn, 2023). Bridging-based ranking has been used to in several high-impact applications such as ranking crowd-sourced fact checks on X, Instagram, Threads, Facebook, and TikTok (Wojcik et al., 2022; Meta, 2025; TikTok, 2025) and selecting comments in collective response systems like Pol.is and Remesh (Small et al., 2021a; Konya et al., 2023a; Huang et al., 2024).
@@
Fig. 1: Our approach compared to the standard approach to ranking comments. In this example, a platform with $n=6$ users wants to select $k=3$ comments to highlight while optimizing a given score function $f$ (e.g. engagement, civility, diverse approval, etc). This example is inspired by our experiments ranking comments related to campus protests in Sec 6. In the standard approach, the platform simply selects the $k$ comments with the highest score (in this case, diverse approval across the red and blue users). However, this leads to only two users having a comment that they approve of in the selected set $S^{*}$. The two remaining red and blue users do not receive representation despite the fact that they each form a group large enough to warrant representation by proportionality (each of size $\lceil n/k\rceil\geq\lceil 6/3\rceil=2$) and are minimally cohesive (have at least one item that they agree upon in common). On the other hand, in our approach, the platform picks the $k$ comments that maximize the score while satisfying justified representation, and guarantees that all cohesive groups of size at least $n/k$ are represented in the resulting set $S^{*}_{JR}$.
@@
While these interventions may improve conversational quality, the goal of ensuring representation in online discussions is less studied. In fact, solely focusing on conversational quality might inadvertently result in the censorship of certain groups, in the sense that their comments may be filtered out or do not appear high enough in the comment ranking to gain visibility. For example, civility classifiers have been shown to disproportionately flag comments written in African-American English (Sap et al., 2019) and comments discussing encounters with racism (Lee et al., 2024). In the context of diverse approval, the groups being bridged have a large effect on what content gets shown. Commonly, these groups correspond to political groups like left- or right-leaning users (Wojcik et al., 2022). Therefore, it is plausible that diverse approval could give more visibility to moderate comments, while failing to provide representation of others who hold more ideologically diverse views. [^1] Or, imagine an online platform popular with fans of the Celtics and Knicks (two U.S. basketball teams) where the only posts that get diverse approval are those about Kadeem Allen (a former player in both franchises). Bridging-based ranking à la Ovadya & Thorburn (2023) may overly represent the Kadeem Allen posts, downgrading legitimate interest of the broader fan-bases.
@@
    Theoretical framework. We contribute a theoretical framework in which the platform selects the top-ranked comments by maximizing a score function $f$ (e.g. civility, bridging objectives) subject to a representation constraint known as justified representation (JR) (Aziz et al., 2017) from the social choice literature on approval voting in multi-winner elections.
@@
    Theoretical analysis. We theoretically analyze how the JR constraint affects the platform’s ability to maximize different classes of score functions. Despite negative results in worst-case settings, we show that in the natural setting where users’ opinions cluster into a few groups, enforcing representation does not come at a price to other objectives such as civility or user engagement (§ 5).
@@
    Method and empirical analysis. Leveraging approximation algorithms from social choice, we implement a JR constraint for ranking comments on Remesh [^2] Remesh (Konya et al., 2023a) is a popular collective response system (Ovadya, 2023) used frequently by governments, non-profits, and corporations to elicit opinions from a collective at scale. related to campus protests. We find that enforcing JR significantly enhances representation without compromising other measures of conversational quality. Specifically, we evaluate the price of JR for metrics beyond simple engagement, including diverse approval and content-based bridging classifiers (§ 6), and find a low price in all settings, as predicted by our theoretical results (§ 5).
@@ 2 Related work
Algorithmic interventions for comment ranking focus primarily on promoting conversational quality  (Kolhatkar & Taboada, 2017; Saltz et al., 2024; Lees et al., 2022; Ovadya & Thorburn, 2023; Jia et al., 2024; Piccardi et al., 2024), but without guarantees on representation of viewpoints—another key ideal of deliberation. To incorporate such guarantees into these methods, we focus on selecting the set of top $k$ comments, which allows us to leverage notions of proportional representation. While other research on diversity in recommender systems also focuses on selecting optimal sets of items, diversity in this context is typically defined based on the content of items, e.g., choosing videos from a variety of topics (Kunaver & Požrl, 2017; Zhao et al., 2025). In contrast, our goal here is to select a set of comments that ensures representation, thereby focusing on the perspective of users rather than the content diversity of the items.
@@
To formalize representation, we borrow the social-choice-theoretic concept of justified representation (JR), developed by Aziz et al. (2017) in the context of approval voting in multi-winner elections. Justified representation is an axiom that formalizes a notion of proportionality: it guarantees that every large enough group of users that has shared preferences (approves of at least one comment in common) is allocated at least one of the top comments. While prior work has focused on selecting or generating comments subject to JR or extensions of JR (Halpern et al., 2023; Fish et al., 2024; Bernreiter et al., 2024), these works do not consider the optimization of other exogenous objectives (e.g. content-based classifiers for civility or other attributes) that are a key component of real-world comment ranking. Therefore, we study a constrained optimization problem where the platform selects the top comments based on optimizing a score function (e.g. civility, diverse approval, engagement, etc) subject to JR.
@@
For the comment ranking setting, in particular, JR affords several advantages over notions of demographic or social representation more commonly used in the algorithmic fairness literature (Barocas et al., 2023; Chasalow & Levy, 2021). We discuss these advantages further in Appendix A, but here, we highlight two benefits. First, unlike most algorithmic fairness methods, JR does not require inference of any demographic or social labels, making it compatible with the privacy and legal constraints of real-world platforms (Holstein et al., 2019; Veale & Binns, 2017). Second, JR automatically adapts to the groups that are relevant in different contexts. For instance, the groups relevant to represent in the discussions about the NYC budget planning process differ from those in a post about a basketball game between the Celtics and the Knicks. On real-world platforms that have a diverse range of content, the fact that JR naturally focuses on the relevant groups is a significant advantage.
The primary theoretical question we analyze is: what is the price of enforcing JR, with respect to the ability to optimize other score functions? In the recommender system context, it is essential that a representation constraint be compatible with other objectives of interest for the platform (e.g. civility of comments). The need to optimize other objectives well is also why we start with the fundamental JR axiom rather than its stronger extensions (Aziz et al., 2017; Sánchez-Fernández et al., 2017; Skowron et al., 2017; Brill & Peters, 2023; Fish et al., 2024; Peters & Skowron, 2020). Nevertheless, in our empirical experiments, we find that all but one of our JR committees also satisfies the stronger EJR+ axiom (Brill & Peters, 2023), consistent with other studies that find minimal differences between these axioms in real-world settings (Boehmer et al., 2024). [^3] See also Bardal et al. (2025) for similar findings beyond the approval voting setting.
@@
The closest related theoretical work is by Elkind et al. (2022, 2024) who investigate the price of JR with respect to utilitarian social welfare (the number of selected comments that each user approves of, summed over users). Other related work includes Skowron et al. (2017) who study the performance of various approval-based committee rules with respect to social welfare and coverage and Bredereck et al. (2019) who establish that maximizing social welfare or coverage with respect to JR is NP-hard. In contrast to Elkind et al. (2022), we analyze the price of JR for more general classes of score functions, which are relevant in the context of recommender systems where objectives beyond user approval are also important, as well as constraints on the user approval matrix. We show that in the common setting when users’ preferences cluster into a few groups, the price of JR, even with respect to arbitrary objective functions, is typically negligible (§ 5). These findings also relate to Faliszewski et al. (2018) who use clustering as an empirical heuristic to achieve fully proportional representation.
@@ 3.1 Model
For $m\in\mathbb{N},$ we write $[m]=\{1,\dots,m\}.$ Let $[m]$ be the set of all comments and $[n]$ be a set of users. The platform must select a subset $S\subseteq[m]$ of $k$ comments to highlight. [^4] Work on diversity in recommender systems also focuses on sets of items, but with the goal of showing “diverse” items where diversity is typically defined based on the content of the items, e.g., showing videos spanning different topics or genres (Kunaver & Požrl, 2017; Zhao et al., 2025). In contrast, our focus is on selecting sets of items (comments) to ensure that the top items provide a degree of proportional representation to users. Each user $u\in[n]$ approves a set $A_{u}\subseteq[m]$ of comments and $\mathcal{A}_{n}=\{A_{1},\dots,A_{n}\}$ is the profile of approved comments across all users. In practice, a user’s approval set might be defined as the set of comments that they upvote, like or react to. In all, each problem is characterized by a tuple $\mathcal{I}_{m,\mathcal{A}_{n},k}=\langle m,\mathcal{A}_{n},k\rangle$ of $n$ approval sets over $m$ comments, from which $k$ comments are selected.
@@ 3.2.2 Approval Dependent Scoring Rules
A scoring rule $f(i,\mathcal{A}_{n})$ is approval dependent if, for all items $i$ not in any approval set of $\mathcal{A}_{n}$ (that is $i\notin\bigcup_{u=1}^{n}A_{u}$), then we have $f(i,\mathcal{A}_{n})=0$.
@@
Diverse approval is another example of an approval dependent scoring rule, where the score of a comment reflects the level of approval it receives across diverse groups of users. These groups could be pre-defined, e.g., based on demographic characteristics, or more commonly, learned from the data. Diverse approval is motivated by research finding that, compared to pure engagement, diverse approval tends to correlate with comments that are less toxic, more informative, and higher quality (Ovadya & Thorburn, 2023; Wojcik et al., 2022). Diverse approval has been used to select and rank user-generated content in high-impact applications such as ranking crowd-sourced fact checks on social media platforms (Wojcik et al., 2022; TikTok, 2025; Meta, 2025), and selecting comments in collective response systems like Pol.is and Remesh (Small et al., 2021a; Konya et al., 2023a; Huang et al., 2024).
@@
Other variants have also been used such as the product of the approval rate across groups (Small et al., 2021a; Huang et al., 2024) or a softmax version of minimax diverse approval (Konya et al., 2023a). Note that these examples satisfy a property that is strictly stronger than approval dependency, that of approval monotonicity where an approval monotonic function is an approval dependent function such that for all approval profiles ${A}^{\prime}_{n}$, and $\mathcal{A}_{n}$ with $A^{\prime}_{u}\succeq_{(i)}A_{u}$ for some $u$ and $A^{\prime}_{v}=A_{v}$ for all $v\in[n]\setminus\{u\},$ then $f(i,\mathcal{A}^{\prime}_{n})\geq f(i,\mathcal{A}_{n}).$
@@ 3.3 Justified Representation
The problem of selecting $k$ items from a set $[m]$ based on individual’s approval ballots is well studied in the context of multi-winner social choice, where the $k$ “items” are candidates to be selected for a committee. [^6] An intuitive approach would be to select the $k$ items with the largest approval score |$|\{u\in[n]\mid i\in A_{u}\}|$. However, this method tends to favor items supported by the majority, potentially excluding minority groups from representation. Let a world with $60$ people who approve the $10$ items in set $A$ and another $40$ people who approve of a distinct set $A^{\prime}$ of $10$ items. If a committee of size $10$ is selected based on approval scores, the winning committee, $A$ would fail to represent $40\%$ of the world. Instead, a committee composed of $6$ items from $A$ and $4$ items from $A^{\prime}$ would respect an intuitive notion of proportionality. In order to ensure fairness, in the sense that all large groups with cohesive preferences receive some amount of representation, (Aziz et al., 2017) axiomatically formalized the idea of justified representation. Following (Elkind et al., 2022)’s exposition, and (Bredereck et al., 2019)’s notion of justifying sets, assume we are given an instance $\mathcal{I}_{m,\mathcal{A}_{n},k}.$ of $n$ approval sets over $m$ comments, from which the platform must select $k$ comments.
@@
The concept of Justified Representation (JR) ensures that large enough cohesive groups are minimally represented.